OpenStudy (anonymous):

Calculate the energy released when 1.00 kilograms of mass is converted entirely to energy. [(1 J = 1 kg times m squared over s squared ) and the speed of light = 3.00 × 108 m over s ] Answer 3.00 × 108 J 9.00 × 1016 J 1.00 × 10-7 J 1.11 × 10-17 J

OpenStudy (abhisar):

LOL i just did it in ur previous question !

OpenStudy (abhisar):

C stands for the speed of light. So for loss of say 1Kg we get an energy equivalent to \(\sf (3\times 10^8)^2=9\times 10^{16}Joules\)

OpenStudy (abhisar):

Got it ?

OpenStudy (abhisar):

That result is obtained using \(\large E=mc^2\)

OpenStudy (anonymous):

so this is the one you were talking about? hahaha

OpenStudy (anonymous):

or are both the answers you gave on both questions correct?

OpenStudy (abhisar):

yes ofcrse...!

OpenStudy (anonymous):

so to the e=mc2 it was c also? and this one is c?

OpenStudy (abhisar):

no for both it's b

OpenStudy (anonymous):

dont do that, its late and im sleep deprived lol

OpenStudy (abhisar):

lol..m serious for both it's b...what are you doing ? i solved the whole question and despite that u r not able to choose the correct option ?

OpenStudy (anonymous):

ok so this answer is b. what other question are you talking about because i think i have three questions open

OpenStudy (abhisar):

What is your other question ?

OpenStudy (anonymous):

What is the binding energy for a C-12 nucleus with a mass defect of 0.0993 amu? [(1 amu= 1.66 × 10-27 ,1 J = 1 kg times m squared over s squared ) and the speed of light = 3.00 × 108 m over s ] Answer 1.48 × 10-11 J 1.24 × 10-12 J 4.12 × 10-21 J 8.94 × 1015 J

OpenStudy (abhisar):

You can solve this similarly using \(\sf E=mc^2\) \(\sf E=0.0993\times 1.66 \times 10^{-27}\times (3\times 10^8)^2\)

OpenStudy (anonymous):

easy way?

OpenStudy (abhisar):

That's the easiest way !

OpenStudy (abhisar):

Come on u just have to multiply the terms...

OpenStudy (kirbykirby):

yep I agree.. lol

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar You can solve this similarly using \(\sf E=mc^2\) \(\sf E=0.0993\times 1.66 \times 10^{-27}\times (3\times 10^8)^2\\ \Rightarrow E=1.48\times 10^{-27}\times 10^{16}\\ \Rightarrow E= ?\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

1.48 X 10 to the negative 11

OpenStudy (kirbykirby):

yeah

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (anonymous):

Which of the following are true for E = mc2? Answer implies that a small change in mass produces a great deal of energy mainly used for chemical reactions implies that a small change is mass produces a small amount of energy

OpenStudy (anonymous):

how do i do the comments like you are doing lol

OpenStudy (abhisar):

Now, observe urself......change in mass of 1kg produces an energy of \(\sf 9\times 10^{16}\). So is it large energy produced or small energy produced ?

OpenStudy (anonymous):

large

OpenStudy (abhisar):

So the correct option should be ?

OpenStudy (anonymous):

for the previous question or the one i just posted here?

OpenStudy (abhisar):

for the one you just posted...this is the reason i ask everyone to post one question per thread.

OpenStudy (anonymous):

sorry i was trying to keep things simple since im always asking questions

OpenStudy (abhisar):

anyways....is every question clear to you ?

OpenStudy (anonymous):

it was 1.48 X 10 to the negative 11

OpenStudy (anonymous):

yes except the one about what is true for the formula e=mc2

OpenStudy (abhisar):

ask it in a new post